# Structural-Spectral Graph Convolution with Evidential Edge Learning for Hyperspectral Image Clustering

### v3

@@ I Introduction
Hyperspectral image (HSI) captures the reflectance of land covers across hundreds of contiguous narrow bands, providing rich and precise spectral information. Classifying the pixels of HSI into different categories is a fundamental task [1, 2, 3, 4, 5, 6, 7], as it describes the spatial distribution of different land covers. This technique is widely applied in various fields, such as precision agriculture [8], environmental monitoring [9], mineral exploration [10], and urban planning [11].
@@
Due to the high cost of obtaining HSI pixel annotations, unsupervised clustering methods that explore the relationships between pixels without labels have gained increasing attention. Traditional clustering methods, e.g., K-means [12], fuzzy c-means [13] and subspace clustering [14], have been applied to HSI clustering, but with relatively poor clustering performance. This is because those methods usually perform clustering based on the raw feature of HSI, which is susceptible to spectral noise of HSI caused by the illumination condition or the viewing angle [15]. To this end, research has increasingly turned to deep clustering methods [16, 17, 18], which leverage deep neural networks to learn discriminative feature representations, achieving effective clustering results under end-to-end training. However, the massive number of pixels in HSI may lead to high computational cost. Consequently, many studies employed superpixel segmentation techniques [19, 20, 21] and transformed the clustering task from the pixel level to the superpixel level to reduce the computational cost. Superpixel-level clustering has currently become the mainstream approach for large-scale HSI clustering [22, 23, 24, 25, 26].
@@
Given the irregular shapes and clear spatial distribution, superpixels are naturally modeled as graph-structured data, where each superpixel acts as a node and the topological graph is constructed based on their spatial adjacency or spectral similarity. With the powerful capability of information aggregation, graph neural networks (GNNs) have excelled in graph learning tasks [27, 28]. Consequently, several studies have employed GNNs for superpixel representation learning to incorporate topological information into superpixel embeddings and then conducted GNN-based clustering, which can be mainly categorized into two types. The first type is based on graph autoencoders (GAEs) [29], which typically reconstruct the graph adjacency matrix from node embeddings, and introduce an auxiliary distribution loss [16] for self-supervised training [30, 31, 32, 33, 34]. The second type utilizes contrastive clustering frameworks [35, 36, 37], conducting alignment or contrast between samples or prototypes to improve the discriminability of superpixel representations [17, 18, 38]. Although the above methods have achieved promising results, two critical limitations still exist.
@@
The first limitation concerns insufficient feature extraction from superpixels. Besides the graph structure containing rich spatial information, the spectra of superpixels exhibit continuity and redundancy [39]. While existing GNNs can effectively aggregate spatial information within the neighborhood, the projective transformations (i.e., linear layers) in their networks struggle to capture complex patterns of HSI spectra. Conversely, convolutional neural networks (CNNs) are adept at extracting high-order spectral features, but fail to exploit the superpixel topological information. To illustrate representation learning abilities of different networks, in Fig. 1a, we conducted a comparison among MLP, 1D convolution (1D-Conv), graph convolution (Graph-Conv) and our proposed operator designed to extract both spatial and spectral features, indicating that comprehensive superpixel feature extraction significantly enhances the clustering accuracy.
@@
The second limitation is the inaccurate superpixel topological graph. Typically, the superpixel topological graph is constructed based on either spatial adjacency or spectral similarity. The former connects spatially adjacent superpixels as neighbors [25, 38], while the latter determines neighbors based on spectral similarity [30, 32, 33]. EGRC-Net[31] adaptively refines the initial graph by adding edges between the nearest node pairs based on nodes’ current embeddings. However, it fails to take into account the reliability of existing edges. Since the superpixel topological graph is constructed without supervision, both spatial and spectral graphs inevitably contain incorrect edges that connect two superpixel nodes of different classes. Accordingly, these incorrect edges bring semantically different neighbors, leading to distorted node representations and confused class semantics during GNN aggregation. Fig. 1b illustrates the impact of deleting incorrect edges on clustering results. As the deletion ratio increases, clustering performance gradually improves, highlighting the importance of dealing with incorrect edges in the graph.
@@ II Data Preprocessing
To reduce the computational cost on large HSI, we employ entropy rate superpixel segmentation (ERS) [19] to segment the input HSI $\mathcal{H}$ into $M$ non-overlapped superpixels, where each pixel belongs to exactly one superpixel. Specifically, we perform principal components analysis (PCA) on pixel features and then retain the first principal component to obtain the gray-scale image as the input of ERS.
Then, we apply PCA on raw pixel features again, and retain only $d$ components to further reduce the data dimensionality [33]. For the $i$-th pixel $p_{i}$ and the $j$-th superpixel $sp_{j}$, their feature vectors are denoted as $\bm{x}^{p}_{i}\in\mathbb{R}^{d}$ and $\bm{x}_{j}\in\mathbb{R}^{d}$, following:
@@
where $\mathrm{dis}(\cdot)$ denotes the Manhattan distance. Following [27], we add self-loop to $\mathbf{A}$ and employ the re-normalization trick:
@@ III-A Contrastive Clustering Framework
We achieve HSI superpixel clustering under a contrastive clustering framework. Specifically, following [36, 37], we adopt an online network $f(\cdot)$ and a target network $f^{\prime}(\cdot)$ with identical architecture but unshared parameters as superpixel encoders. To break the symmetry between $f(\cdot)$ and $f^{\prime}(\cdot)$, and prevent representation collapse, predictor $g(\cdot)$ consisting of two-layer MLP is introduced after $f(\cdot)$. Neighborhood alignment and prototype contrast are combined as clustering-oriented training objectives.
For neighborhood alignment, we first construct reliable positive samples for superpixels through data augmentation. Given superpixel $\bm{x}_{i}$, we randomly select one internal pixel as its augmented view $\bm{x}_{i}^{+}$ [38]. Then, based on the assumption that representations within a local neighborhood in the embedding space share similar semantics, superpixel representations are encouraged to be similar to not only their augmented views but also neighboring samples, thus improving intra-cluster compactness. Instead of directly sampling from the neighborhood, adopting the reparameterization trick ensures proper gradient back-propagation. For superpixel $\bm{x}_{i}$, its neighborhood in the embedding space can be represented as:
@@
Given one prototype, the remaining K1K-1 prototypes are definitely negative examples. We construct a contrastive loss [40] to encourage the alignment between different views of the same prototype, and increase the inter-cluster distance, i.e.,
@@ III-B Structural-Spectral Graph Convolutional Operator
where $\mathbf{H}_{(l)}$ is the output of the $l$-th layer and $\mathbf{H}_{(0)}=\mathbf{X}$, $\circledast_{\scriptscriptstyle\mathrm{1D}}$ denotes 1D convolutional operator, $\mathbf{W}^{c}_{(l)}$ and $\mathbf{W}^{g}_{(l)}$ are trainable parameters in the $l$-th layer corresponding to 1D convolution and graph convolution, $\phi(\cdot)$ is non-linear activation function. Moreover, we conduct batch normalization (BN) [41] after each layer to stabilize the training.
@@ IV-B Experiment Setup
1) Implementation Details: The input HSI is standardized across each spectral band. To decrease computational costs, PCA is adopted to reduce the number of bands to 40, 20, 25 and 40 on Indian Pines, Pavia University, Botswana and Trento, respectively. In the contrastive clustering framework, the hidden size of the predictor $g(\cdot)$ is set to 512. The momentum coefficient $m$ for updating $f^{\prime}(\cdot)$ is set to 0.99. For neighborhood alignment, $\sigma$ is set to $1\times 10^{-3}$. For prototype contrast, $\tau$ is set to 0.7. In the structural-spectral graph convolutional operator, the kernel size of 1D convolution in the first layer is set to 7, with subsequent layers decreasing by 2, but not smaller than 3. The output channels in the first layer is set to 8, with subsequent layers multiplying by 2, but not larger than 64. FC layers in graph convolution keep output dimensions equal to input dimensions. In the evidence-guided adaptive edge learning, $h(\cdot)$ consists of a two-layer MLP with the hidden size of $K$. We used the SGD optimizer with the learning rate, weight decay, and momentum coefficient set to 0.05, $5\times 10^{-4}$, and 0.9, respectively. In addition, we adopted cosine annealing and set the learning rate of $g(\cdot)$ $10\times$ as that of other modules following [36]. We repeated all experiments five times and reported the mean and standard deviation of the results. We carried out all experiments in a Linux server with an AMD EPYC 7352 CPU and an NVIDIA RTX 4090 GPU.
@@
2) Compared Methods: To thoroughly assess the clustering performance of the proposed method, we compare it against eleven baseline methods, including K-means [12], SSC [14], NCSC [26], SGLSC [25], DGAE [33], EGRC [31], S3-ULDA [42], S2DL [43], SDST [32], SPGCC [38] and SAPC [44]. For K-means, clustering is conducted at the pixel level. In contrast, SSC performs clustering at the superpixel level to mitigate the high memory cost. For other methods, we follow the settings in their papers and exhaustively search optimal values for hyperparameters. Here we briefly introduce each method:
@@
K-means [12] is a traditional centroid-based clustering algorithm with “K-means++” initialization.
@@
SSC [14] is a traditional subspace-based clustering algorithm with the sparse constraint and considering the outliers.
@@
NCSC [26] introduces neighborhood contrastive regularization into deep subspace clustering.
@@
SGLSC [25] constructs global and local similarity graphs and performs subspace clustering.
@@
DGAE [33] builds superpixel spatial and spectral graph and proposes a dual graph autoencoder.
@@
EGRC [31] refines the initial graph by adding edges between the nearest node pairs based on nodes’ current embeddings and is trained under the autoencoder framework.
@@
S3-ULDA [42] is a superpixel dimensionality reduction method based on LDA, followed by K-means.
@@
S2DL [43] performs diffusion learning based on superpixels and spatial regularization to achieve geometry-based clustering.
@@
SDST [32] proposes a double-structure transformer with shared autoformer and achieves clustering under the autoencoder framework.
@@
SPGCC [38] proposes superpixel graph contrastive clustering with semantic-invariant augmentations.
@@
SAPC [44] constructs pixel-anchor and anchor-anchor graphs and proposes structured anchor projected clustering.
@@
3) Evaluation Metrics: To quantitatively evaluate the clustering performance of different methods, we adopt eight mainstream clustering metrics: accuracy (ACC), kappa coefficient ($\mathcal{K}$), normalized mutual information (NMI), adjusted Rand index (ARI), Precision, Recall, F1-score (F1) and Purity. For Precision, Recall and F1, we use the macro-average form. Among these metrics, ACC, NMI, Precision, Recall, F1 and Purity range in [0, 1] while Kappa and ARI range in [-1, 1]. Each metric is reported as a percentage, with a higher value indicating the better clustering result. Following the common process, we adopt the Hungarian algorithm [45] to map the predicted labels based on the ground truth before calculating clustering metrics.
@@ IV-C Comparison with State-of-the-Art Methods
According to the quantitative clustering metrics, it could be observed that our method outperforms other compared methods with a large margin. For instance, our method improves the ACC by 2.61%, 6.06%, 4.96% and 3.15% on the four datasets. Notably, on Pavia University, ACC, $\mathcal{K}$, ARI and F1 obtained by our method outperform the second-best value by 6.06%, 7.47%, 8.22% and 1.13%, respectively. Overall, on Botswana and Trento, our method achieves the best scores on all metrics; on Indian Pines, it achieves the best scores on five metrics and the second-best on three metrics; on Pavia University, it achieves the best scores on seven metrics and the second-best on one metric. Based on the results of pairwise $t$-Test at 0.05 significance level, our method significantly outperforms the compared methods in 91.8% (323/352) cases.
@@ IV-E Running Time Analysis
We report the running time of our method and other compared methods on four datasets in Table IV. In terms of the average running time, our method is only slower than SSC and SAPC, while faster than other nine methods, demonstrating that our method is lightweight and efficient. Compared to SSC, our method achieves significantly better clustering performance on all datasets. Compared to SAPC, our method demonstrates an improvement of over 10% in ACC on Indian Pines and Botswana, and an improvement of 4%-6% on other datasets. Therefore, the slight increase in running time is acceptable.
@@
Network variations     Indian Pines     Pavia University     Botswana     Trento
@@
Items     Indian Pines     Pavia University     Botswana     Trento

### v4

@@ I Introduction
Hyperspectral image (HSI) captures the reflectance of land covers across hundreds of contiguous narrow bands, providing rich and precise spectral information. Classifying the pixels of HSI into different categories is a fundamental task [hsi_cls_1, hsi_cls_2, hsi_cls_3, hsi_cls_4, hsi_cls_5, hsi_cls_6, hsi_cls_7], as it describes the spatial distribution of different land covers. This technique is widely applied in various fields, such as precision agriculture [app_precision_agriculture], environmental monitoring [app_environmental_monitoring], mineral exploration [app_mineral_exploration], and urban planning [app_urban_planning].
@@
Due to the high cost of obtaining HSI pixel annotations, unsupervised clustering methods that explore the relationships between pixels without labels have gained increasing attention. Traditional clustering methods, e.g., K-means [kmeans], fuzzy c-means [fuzzycmeans] and subspace clustering [ssc], have been applied to HSI clustering, but with relatively poor clustering performance. This is because those methods usually perform clustering based on the raw feature of HSI, which is susceptible to spectral noise of HSI caused by the illumination condition or the viewing angle [spectral_noise]. To this end, research has increasingly turned to deep clustering methods [dec, pscpc, s2gcl], which leverage deep neural networks to learn discriminative feature representations, achieving effective clustering results under end-to-end training. However, the massive number of pixels in HSI may lead to high computational cost. Consequently, many studies employed superpixel segmentation techniques [ers, slic, lsc] and transformed the clustering task from the pixel level to the superpixel level to reduce the computational cost. Superpixel-level clustering has currently become the mainstream approach for large-scale HSI clustering [superpixel_clustering_1, superpixel_clustering_2, superpixel_clustering_3, sglsc, ncsc].
@@
Given the irregular shapes and clear spatial distribution, superpixels are naturally modeled as graph-structured data, where each superpixel acts as a node and the topological graph is constructed based on their spatial adjacency or spectral similarity. With the powerful capability of information aggregation, graph neural networks (GNNs) have excelled in graph learning tasks [gcn, gat]. Consequently, several studies have employed GNNs for superpixel representation learning to incorporate topological information into superpixel embeddings and then conducted GNN-based clustering, which can be mainly categorized into two types. The first type is based on graph autoencoders (GAEs) [gae], which typically reconstruct the graph adjacency matrix from node embeddings, and introduce an auxiliary distribution loss [dec] for self-supervised training [lowpass, EGRC, sdst, dgae, graph_clustering]. The second type utilizes contrastive clustering frameworks [moco, byol, propos], conducting alignment or contrast between samples or prototypes to improve the discriminability of superpixel representations [pscpc, s2gcl, spgcc]. Although the above methods have achieved promising results, two critical limitations still exist.
@@
The first limitation concerns insufficient feature extraction from superpixels. Besides the graph structure containing rich spatial information, the spectra of superpixels exhibit continuity and redundancy [survey]. While existing GNNs can effectively aggregate spatial information within the neighborhood, the projective transformations (i.e., linear layers) in their networks struggle to capture complex patterns of HSI spectra. Conversely, convolutional neural networks (CNNs) are adept at extracting high-order spectral features, but fail to exploit the superpixel topological information. To illustrate representation learning abilities of different networks, in Fig. 1a, we conducted a comparison among MLP, 1D convolution (1D-Conv), graph convolution (Graph-Conv) and our proposed operator designed to extract both spatial and spectral features, indicating that comprehensive superpixel feature extraction significantly enhances the clustering accuracy.
@@
The second limitation is the inaccurate superpixel topological graph. Typically, the superpixel topological graph is constructed based on either spatial adjacency or spectral similarity. The former connects spatially adjacent superpixels as neighbors [sglsc, spgcc], while the latter determines neighbors based on spectral similarity [lowpass, sdst, dgae]. EGRC-Net[EGRC] adaptively improves the initial graph by adding edges between the nearest node pairs based on nodes’ current embeddings. However, it fails to take into account the reliability of existing edges. Since the superpixel topological graph is constructed without supervision, both spatial and spectral graphs inevitably contain incorrect edges that connect two superpixel nodes of different classes. Accordingly, these incorrect edges bring semantically different neighbors, leading to distorted node representations and confused class semantics during GNN aggregation. Fig. 1b illustrates the impact of deleting incorrect edges on clustering results. As the deletion ratio increases, clustering performance gradually improves, highlighting the importance of dealing with incorrect edges in the graph.
@@ II Data Preprocessing
To reduce the computational cost on large HSI, we employ entropy rate superpixel segmentation (ERS) [ers] to segment the input HSI $\mathcal{H}$ into $M$ non-overlapped superpixels, where each pixel belongs to exactly one superpixel. Specifically, we perform principal components analysis (PCA) on pixel features and then retain the first principal component to obtain the gray-scale image as the input of ERS.
Then, we apply PCA on raw pixel features again, and retain only $d$ components to further reduce the data dimensionality [dgae]. For the $i$-th pixel $p_{i}$ and the $j$-th superpixel $sp_{j}$, their feature vectors are denoted as $\bm{x}^{p}_{i}\in\mathbb{R}^{d}$ and $\bm{x}_{j}\in\mathbb{R}^{d}$, following:
@@
where $\mathrm{dis}(\cdot)$ denotes the Manhattan distance. Following [gcn], we add self-loop to $\mathbf{A}$ and employ the re-normalization trick:
@@ III-A Contrastive Clustering Framework
We achieve HSI superpixel clustering under a contrastive clustering framework. Specifically, following [byol, propos], we adopt an online network $f(\cdot)$ and a target network $f^{\prime}(\cdot)$ with identical architecture but unshared parameters as superpixel encoders. To break the symmetry between $f(\cdot)$ and $f^{\prime}(\cdot)$, and prevent representation collapse, predictor $g(\cdot)$ consisting of two-layer MLP is introduced after $f(\cdot)$. Neighborhood alignment and prototype contrast are combined as clustering-oriented training objectives.
For neighborhood alignment, we first construct reliable positive samples for superpixels through data augmentation. Given superpixel $\bm{x}_{i}$, we randomly select one internal pixel as its augmented view $\bm{x}_{i}^{+}$ [spgcc]. Then, based on the assumption that representations within a local neighborhood in the embedding space share similar semantics, superpixel representations are encouraged to be similar to not only their augmented views but also neighboring samples, thus improving intra-cluster compactness. Instead of directly sampling from the neighborhood, adopting the reparameterization trick ensures proper gradient back-propagation. For superpixel $\bm{x}_{i}$, its neighborhood in the embedding space can be represented as:
@@
Given one prototype, the remaining K1K-1 prototypes are definitely negative examples. We construct a contrastive loss [infonce] to encourage the alignment between different views of the same prototype, and increase the inter-cluster distance, i.e.,
@@ III-B Structural-Spectral Graph Convolutional Operator
where $\mathbf{H}_{(l)}$ is the output of the $l$-th layer and $\mathbf{H}_{(0)}=\mathbf{X}$, $\circledast_{\scriptscriptstyle\mathrm{1D}}$ denotes 1D convolutional operator, $\mathbf{W}^{c}_{(l)}$ and $\mathbf{W}^{g}_{(l)}$ are trainable parameters in the $l$-th layer corresponding to 1D convolution and graph convolution, $\phi(\cdot)$ is non-linear activation function. Moreover, we conduct batch normalization (BN) [bn] after each layer to stabilize the training.
@@ IV-B Experiment Setup
1) Implementation Details: The input HSI is standardized across each spectral band. To decrease computational costs, PCA is adopted to reduce the number of bands to 40, 20, 25 and 40 on Indian Pines, Pavia University, Botswana and Trento, respectively. In the contrastive clustering framework, the hidden size of the predictor $g(\cdot)$ is set to 512. The momentum coefficient $m$ for updating $f^{\prime}(\cdot)$ is set to 0.99. For neighborhood alignment, $\sigma$ is set to $1\times 10^{-3}$. For prototype contrast, $\tau$ is set to 0.7. In the structural-spectral graph convolutional operator, the kernel size of 1D convolution in the first layer is set to 7, with subsequent layers decreasing by 2, but not smaller than 3. The output channels in the first layer is set to 8, with subsequent layers multiplying by 2, but not larger than 64. FC layers in graph convolution keep output dimensions equal to input dimensions. In the evidence-guided adaptive edge learning, $h(\cdot)$ consists of a two-layer MLP with the hidden size of $K$. We used the SGD optimizer with the learning rate, weight decay, and momentum coefficient set to 0.05, $5\times 10^{-4}$, and 0.9, respectively. In addition, we adopted cosine annealing and set the learning rate of $g(\cdot)$ $10\times$ as that of other modules following [byol]. We repeated all experiments five times and reported the mean and standard deviation of the results. We carried out all experiments in a Linux server with an AMD EPYC 7352 CPU and an NVIDIA RTX 4090 GPU.
@@
2) Compared Methods: To thoroughly assess the clustering performance of the proposed method, we compare it against eleven baseline methods, including K-means [kmeans], SSC [ssc], NCSC [ncsc], SGLSC [sglsc], DGAE [dgae], EGRC [EGRC], S3-ULDA [s3ulda], S2DL [s2dl], SDST [sdst], SPGCC [spgcc] and SAPC [sapc]. For K-means, clustering is conducted at the pixel level. In contrast, SSC performs clustering at the superpixel level to mitigate the high memory cost. For other methods, we follow the settings in their papers and exhaustively search optimal values for hyperparameters. Here we briefly introduce each method:
@@
K-means [kmeans] is a traditional centroid-based clustering algorithm with “K-means++” initialization.
@@
SSC [ssc] is a traditional subspace-based clustering algorithm with the sparse constraint and considering the outliers.
@@
NCSC [ncsc] introduces neighborhood contrastive regularization into deep subspace clustering.
@@
SGLSC [sglsc] constructs global and local similarity graphs and performs subspace clustering.
@@
DGAE [dgae] builds superpixel spatial and spectral graph and proposes a dual graph autoencoder.
@@
EGRC [EGRC] refines the initial graph by adding edges between the nearest node pairs based on nodes’ current embeddings and is trained under the autoencoder framework.
@@
S3-ULDA [s3ulda] is a superpixel dimensionality reduction method based on LDA, followed by K-means.
@@
S2DL [s2dl] performs diffusion learning based on superpixels and spatial regularization to achieve geometry-based clustering.
@@
SDST [sdst] proposes a double-structure transformer with shared autoformer and achieves clustering under the autoencoder framework.
@@
SPGCC [spgcc] proposes superpixel graph contrastive clustering with semantic-invariant augmentations.
@@
SAPC [sapc] constructs pixel-anchor and anchor-anchor graphs and proposes structured anchor projected clustering.
@@
3) Evaluation Metrics: To quantitatively evaluate the clustering performance of different methods, we adopt eight mainstream clustering metrics: accuracy (ACC), kappa coefficient ($\mathcal{K}$), normalized mutual information (NMI), adjusted Rand index (ARI), Precision, Recall, F1-score (F1) and Purity. For Precision, Recall and F1, we use the macro-average form. Among these metrics, ACC, NMI, Precision, Recall, F1 and Purity range in [0, 1] while Kappa and ARI range in [-1, 1]. Each metric is reported as a percentage, with a higher value indicating the better clustering result. Following the common process, we adopt the Hungarian algorithm [hungarian] to map the predicted labels based on the ground truth before calculating clustering metrics.
@@ IV-C Comparison with State-of-the-Art Methods
According to the quantitative clustering metrics, it could be observed that our method outperforms other compared methods by a large margin. For instance, our method improves the ACC by 2.61%, 6.06%, 4.96% and 3.15% on the four datasets. Notably, on Pavia University, ACC, $\mathcal{K}$, ARI and F1 obtained by our method outperform the second-best value by 6.06%, 7.47%, 8.22% and 1.13%, respectively. Overall, on Botswana and Trento, our method achieves the best scores on all metrics; on Indian Pines, it achieves the best scores on five metrics and the second-best on three metrics; on Pavia University, it achieves the best scores on seven metrics and the second-best on one metric. Based on the results of pairwise $t$-Test at 0.05 significance level, our method significantly outperforms the compared methods in 91.8% (323/352) cases.
@@ IV-E Running Time Analysis
We report the running time of our method and other compared methods on four datasets in Table IV. In terms of the average running time, our method is only slower than SSC and SAPC, while faster than the other nine methods, demonstrating that our method is lightweight and efficient. Compared to SSC, our method achieves significantly better clustering performance on all datasets. Compared to SAPC, our method demonstrates an improvement of over 10% in ACC on Indian Pines and Botswana, and an improvement of 4%-6% on other datasets. Therefore, the slight increase in running time is acceptable.
@@
Network Variation     Indian Pines     Pavia University     Botswana     Trento
@@
Item     Indian Pines     Pavia University     Botswana     Trento